OpenStudy (anonymous):

Which of the formula(s) is most likely to help you relate side(s) to an angle? -Law of Cosines. -K=1/2ab sinC. -Law of Sines. -Hero's Formula.

OpenStudy (anonymous):

Law of sines

OpenStudy (anonymous):

Law of cosine !?

OpenStudy (anonymous):

what is the difference ?

OpenStudy (anonymous):

Does anyone know about the second one maybe it's from property of triangles ?

OpenStudy (anonymous):

if you know what those laws are, you will easily understand the difference.

OpenStudy (anonymous):

Both Sine and Cosine laws

OpenStudy (anonymous):

yes sines and cosines laws

OpenStudy (anonymous):

Chenna is such an retrice lol

OpenStudy (anonymous):

if you dont mind your language you ll get your retricekicked out of this earth.

OpenStudy (zarkon):

that would be a really hard kick ;)

OpenStudy (anonymous):

what? o.O

OpenStudy (anonymous):

Given angles: phi, theta, rho. Given sides: alpha,beta,gamma (opposite respective angles) The law of sines: \[\frac{\sin(\phi)}{\alpha}=\frac{\sin(\theta)}{\beta}=\frac{\sin(\rho)}{\gamma}\] The law of cosines is: \[\gamma^2=\alpha^2+\beta^2-2\alpha \beta \cos(\rho)\] I was just saying, you don't have to have such a condescending, "why are you so ignorant" tone.

OpenStudy (anonymous):

@ishaan the second one is the formula for the area of a triangle. \[area = \frac{1}{2}absin(C)\] works if you have two sides and the angle in between them.

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

its really waste of time to want to help you guys with people like 'malevolence19' around here. first learn manners and respecting others, then learn maths or physics or any other subject.

OpenStudy (anonymous):

You're level 11. I've obviously helped more people than you. With your little snarky remarks, why don't you take yourself to the MIT groups. I'm sure you'll get alone with them fine.

OpenStudy (anonymous):

joemath the question amistre posted is there something creepy about it I don't know.. you might want to check it

OpenStudy (anonymous):

sure thing~

OpenStudy (anonymous):

Just because you helped more people does not make you any great. As long you don't respect others others wont respect you. So first learn how to behave before you teach others anything. And you might have helped more because you are here for a long time. I had joined this group just yesterday.

OpenStudy (anonymous):

I've been here for about 2 months. 2 weeks of that being away from my PC. So 6 weeks. I respect others, you talk down to them.

OpenStudy (anonymous):

All malevolence is trying to say is that your tone was a little condescending. While things my be "easily seen" or "trivial" to you, they are not to the person asking the question. It might be better if you leave those words out of your responses.

OpenStudy (anonymous):

Precisely.

OpenStudy (anonymous):

chenna what malevolence said was quite right...I guess you are south Indian..and most probably new to all this ..(No offence ) but you are interpreting it wrong. Just leave this.